OpenStudy (anonymous):

An apparel shop sells skirts for $45 and blouses for $35. Its entire stock is worth $51,750, but sales are slow and only half the skirts and two-thirds of the blouses are sold, for a total of $30,600. How many skirts and blouses are left in the store?

OpenStudy (anonymous):

Let's call b for blouses and s for skirts. Set up two equations like this: \[45s+35b=51,750\] \[\frac{ 1 }{ 2 }45s+\frac{ 2 }{ 3 }35b=30,600\] I'm assuming you do know how to solve two linear equations. My values came out to be: \[skirts=520\] \[blouses=810\] Now, these are the "original" number of skirts and blouses, i.e., before the sales. We are asked to find out how many "remain" after the sales. We are told that only half of the skirts were sold: \[skirts-after-the-sales = \frac{1}{2}520=260\] \[blouses-after-the-sales = \frac{1}{3}810=270\]

OpenStudy (anonymous):

can u please show me how u got the answer

OpenStudy (anonymous):

thus the 520 nd 810

OpenStudy (anonymous):

oh ok.let me work it out now i fin i got the idea